OpenStudy (anonymous):

Given the sequence 7, 14, 28, 56, ..., which expression shown would give the tenth term?

OpenStudy (anonymous):

L=A*R^(N-1) L is the last number A is the first number R is the ratio N is the nth term -------------------- L=7*2^(10-1) L=7*2^9 L=7*512 L=3,584 is the 10th term in the sequence